• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name sawatzky

    Log entry time 08:38:07 on March 09, 2005

    Entry number 140955

    keyword=Collimator control on LHRS

    Remote control of the LHRS collimator was not working as of 0400 during owl shift. The on-call tech was called and we learned that we had tried pretty much everything that can be done without an access. We decided to take data with the collimator in the current '6 msr' position (position: 0.6167 m) instead of the sieve.

    I spoke with Ed Folts at 0800 and learned that this is a known problem (weak motor, sticky collimator, or something to that effect). He will have the sieve installed manually during controlled access this morning.

    NOTE: We can't count on being able to change the collimator without an access until this is fixed.